2. Two random samples are chosen from the corresponding populations. Assume that these two samples are totally independent. The measurements obtained are as fol- lows: Sample 1 10.6 10.2 10.5 10.3 10.8 9.8 10.6 10.7 10.2 10.0 Sample 2 9.9 9.8 9.6 10.1 10.2 10.1 9.7 9.5 9.6 9.8 Perform a test to determine whether the population means are significantly different from each other at α-.05. (a) Perform a test for the homogeneity of variance between the two groups. Please set up the appropriate Ho and Ha, calculate the test statistic, find P-value or rejection and state your conclusion. (b) Based on the results from part (a), perform a two sample t-test for the equality of the means (versus the two-sided alternative. Please follow similar steps as in
Answer #1

a) To test 2 against H_1:sigma_1^2 eq sigma_2^2

Here

ซึ่ = 0.105, ni-10, s = 0.0579. n2 10

The test statistic can be written as

2122 S-S which under H0 follows a F distribution with n 1, n2-1) df.

We reject H0 at 5% level of significance if F_{obs}>F_{0.05,n_1-1,n_2-1} or P-value < 0.05

Now,

The value of the test statistic ,-1.81347

and critical value 0.025.9.9 4.02599

and P-value = P(Fn-1,n2-1 > Fobs) P(F9,9>1.81347)1 0.805753 0.194247

Since Fob,-1.81347 ג F0025.9.9-4.02599 and P-value < 0.05,  so we fail to reject H0 at 5% level of significance and we can conclude that the population variances are almost equal.

b) Here,

ή-10.37, si-0. 105, n,-10.T2-9.83. s = 0.0579. n2 = 10

The test statistic can be written as

f_(71- P n1 which under H0 follows a t distribution with n_1 + n_2 - 2 df.

We reject H0 at 5% level of significance if P-value < 0.05

where pooled variance s_p^2 = rac{(n_1-1)s_1^2+(n_2-1)s_2^2}{n_1+n_2-2}

Now,

Pooled variance 20.08145

The value of the test statistic tobs = 4.23090

and P-value =

Pltnitn2-2! > tobs) = Plt181 > 4.23090) = 2 * P(t18 > 4.23090) 1-0.999749 0.000251

Since P-value < 0.05, so we reject H0 at 5% level of significance and we can conclude that the two means are significantly different.